# Self-Consistent Reasoning-based Aspect-Sentiment Quad Prediction with Extract-Then-Assign Strategy

### v1

@@ Abstract
In the task of aspect sentiment quad prediction (ASQP), generative methods for predicting sentiment quads have shown promising results. However, they still suffer from imprecise predictions and limited interpretability, caused by data scarcity and inadequate modeling of the quadruplet composition process. In this paper, we propose Self-Consistent Reasoning-based Aspect-sentiment quadruple Prediction (SCRAP), optimizing its model to generate reasonings and the corresponding sentiment quadruplets in sequence. SCRAP adopts the Extract-Then-Assign reasoning strategy, which closely mimics human cognition. In the end, SCRAP significantly improves the model’s ability to handle complex reasoning tasks and correctly predict quadruplets through consistency voting, resulting in enhanced interpretability and accuracy in ASQP.
11footnotetext: Equal Contribution
@@ 1 Introduction
Aspect-based sentiment analysis (ABSA) refers to the task of identifying entity aspects and their associated sentiments (Pontiki et al., 2014). Among various ABSA tasks, the challenging task of predicting quadruplets, including aspect sentiment quad prediction (ASQP) (Zhang et al., 2021a) and aspect-category-opinion-sentiment (ACOS) (Cai et al., 2021), has garnered significant interest in that it can offer comprehensive aspect-level analysis. Specifically, a quadruplet consists of four sentiment elements: aspect term (at), opinion term (ot), aspect category (ac) and sentiment polarity (sp). Recent studies have developed powerful generative methods by fine-tuning language models (LMs) to sequentially generate sentiment quads (Zhang et al., 2021a; Hu et al., 2022; Gou et al., 2023).
@@
To tackle this challenge in ASQP, our approach introduces a two-step reasoning strategy, termed Extract-Then-Assign. This reasoning strategy initially extracts all aspect and opinion terms from an input sentence; subsequently, it assigns each aspect-opinion pair both an aspect category and a sentiment polarity, utilizing predefined sets of categories and polarities. In Figure 1 Right, explicit modeling of the two steps within the ASQP reasoning can elucidate the generation of complete quadruplets. Note that this reasoning process closely mimics human cognition, thereby enhancing the accuracy and interpretability of quad predictions.
In this work, we propose a novel framework for Self-Consistent Reasoning-based Aspect-sentiment quad Prediction, named SCRAP, which predicts aspect-sentiment quads based on the Extract-Then-Assign reasoning (Figure 2). The key idea is to distill the plausible reasoning ability from large language models (LLMs) into our ASQP model. To this end, we first collect diverse reasoning paths via Chain-of-Thought (CoT) prompting on LLMs Wei et al. (2022), and then optimize our model to generate the reasoning followed by quad prediction. Furthermore, SCRAP aggregates its diverse reasoning outputs based on the self-consistency (Wang et al., 2022). This allows for filtering out noisy outputs and achieving more accurate quad predictions.
@@ 2.1 Problem Formulation
Given an input sentence, aspect sentiment quad prediction (ASQP) aims to predict all aspect-sentiment quads $\{(at,ot,ac,sp)\}$. The aspect term $at$ and opinion term $ot$ are detected within the sentence. The aspect category $ac$ and sentiment polarity $sp$ are classified within their respective predefined sets. Please refer to Appendix A for the details.
@@ 2.2 ASQP Reasoning Generation
Mimicking the human cognition process, which first identifies terms and then infers their relations and semantics, we devise Extract-Then-Assign reasoning strategy for ASQP. From an input sentence, it first extracts at and ot pairs, and subsequently infers the corresponding ac and sp by assigning them to elements within predefined sets of categories and polarities.
@@
We generate diverse reasoning paths using LLM based on the proposed reasoning strategy. We employ a few-shot Chain-of-Thought (CoT) prompting (Wei et al., 2022). Formally, given a sentence and its quadruplets, we generate $N$ reasoning paths $\mathcal{R}=[r_{1},r_{2},...,r_{N}]$ that explain how to reach the quadruplets from the sentence. Our prompt is designed to induce the Extract-Then-Assign process, facilitating the generation of plausible reasoning from the LLM. By leveraging the generated reasoning, we seek to provide rationales that guide and explain the ASQP task to our model, enhancing both accuracy and interpretability. With the generated reasoning, we can provide rationales that guide and explain the ASQP task to our model. The prompt can be found in the Appendix B.
@@ 3.1 Experimental Settings
Dataset and Evaluation Metric
We use two datasets, i.e., Rest15 and Rest16, widely used for ASQP task (Zhang et al., 2021a). For reasoning generation (Sec.2.3), we use ChatGPT (gpt-3.5-turbo-16k). [^1] https://chat.openai.com/ As the evaluation metric, we mainly employ the F1 score with the precision (Pre) and recall (Rec). A predicted quad is considered as correct if and only if its all elements are exactly the same as the ground-truth ones.
@@
We compare SCRAP with two discriminative methods: TAS-BERT Wan et al. (2020), Extract-Classify Cai et al. (2021), and four competitive generative methods: GAS Zhang et al. (2021b), Paraphrase Zhang et al. (2021a), DLO Hu et al. (2022), and MvP Gou et al. (2023). For generative methods, we adopt T5-Base and T5-3B as backbone models. For SCRAP, we set $N=16$ and $P=5$ in common, $K=20$ (T5-Base) and $K=25$ (T5-3B) for Rest15, $K=15$ for Rest16. Refer to Appendix D for implementation details.
@@ 3.2 Results and Discussion
Table 1 and Figure 3 present the ASQP performance of various methods. Overall, SCRAP achieves competitive performance, and outperforms the baseline methods for large backbone models. Specifically, in Figure 3, SCRAP shows higher effectiveness when applied to a larger model, exhibiting the largest performance gap between T5-Base and T5-3B. A larger model generally has higher reasoning capability (Li et al., 2023), making it more suitable for SCRAP which leverages reasoning for ASQP. Furthermore, SCRAP generally exhibits high precision, as it filters out inconsistent predictions through consistency voting. Lastly, we analyze the prediction errors of SCRAP and the best competitor, MvP. In Figure 4 Left, the overall error rate of SCRAP is notably lower than MvP.
@@
We investigate the performance of SCRAP with varying $N$ and $K$, which control the number of reasoning paths used for the training and inference, respectively. In Figure 5, we observe that both methods generally achieve higher performance with a greater number of reasoning paths, and the best performance is achieved by leveraging multiple paths for both training and inference. These results show that diverse reasoning for prediction is indeed beneficial in more accurate quad predictions.
@@
Figure 4 Mid and Right provide the case study. In Example 1, the model predicts the correct quad, and it is possible to interpret the process through which the correct answer was reached. Moreover, even if the model fails to predict the correct quad, we can still understand how the prediction failed based on the Extract-Then-Assign reasoning process. In Example 2, we can see that the model fails to extract the correct opinion term, and this leads to inaccurate polarity prediction. This interpretability helps to better understand the model behavior, which is an important strength of SCRAP.
@@ 4 Related Work
Many existing studies have focused on discriminative methods. Early works tried jointly detecting target-aspect-sentiment (Wan et al., 2020) or conducting ACOS with two-stage pipelines (Cai et al., 2020). Beginning with Zhang et al. (2021b), recent studies have started using generative methods. Zhang et al. (2021a) transforms outputs into natural language sentences, Hu et al. (2022) introduces data augmentation that uses various permutations of quad elements, and Gou et al. (2023) aligns training and inference with multi-view prompting.
@@
CoT prompting has shown high effectiveness in inducing models to generate reasoning before reaching an answer Wei et al. (2022); Kojima et al. (2022); Wang et al. (2022). Recent work has focused on distilling the reasoning capabilities of LLMs to smaller LMs Ho et al. (2022); Magister et al. (2022); Hsieh et al. (2023). They elicit rationales for the predictions from LLMs and utilize them to train the LMs, effectively improving their performance.
@@ 5 Conclusion
This paper aims to enhance quad prediction in ASQP using Extract-Then-Assign, which is two-step reasoning strategy. To this end, we propose a SCRAP framework, which generates diverse predictions utilizing the Extract-Then-Assign and selects the final answer by filtering the inconsistent answers through consistency voting. Our framework is the first method to integrate reasoning into the ABSA task, not only achieving state-of-the-art performance, but also significantly enhancing the interpretability of the outputs produced by the model. This confirms its efficacy in predicting quadruplets through reasoning.
@@ 6 Limitations
Despite achieving state-of-the-art performance, our study has three limitations. Firstly, our current reasoning structure, designed to mimic human cognition, may not be the most advanced or effective for the ASQP task. There could exist a more sophisticated reasoning structure that could further enhance the performance of the task. Secondly, our approach is influenced by the size of the model. With small models having limited reasoning capabilities, our method may not exhibit satisfactory performance. Lastly, our study incurs higher computational costs for training and inference compared to previous studies, as it additionally leverages reasoning.
@@ 7 Ethical Statment
We utilized datasets that are widely recognized and previously employed in the scientific community, maintaining transparency and integrity in our experiments. Our methodologies and findings do not inflict harm upon any individuals or groups. We are cognizant of the potential biases in sentiment polarity predictions arising from the use of large pre-trained language models, as these models may mirror existing societal biases found in their training corpora (Tan and Celis, 2019). We acknowledge the importance of ongoing efforts to mitigate such biases. Furthermore, we underscore the necessity of continuous monitoring and evaluation to ensure that our smaller downstream models do not replicate or amplify the biases inherent in their larger language model counterparts.
@@ Appendix C Target Construction Details
We construct the prediction targets by combining the generated reasoning with the ground-truth quadruplets. To represent the quadruplets, we use special markers: [AT], [OT], [AC], [SP], which respectively denote at, ot, ac, sp, as done in (Hu et al., 2022). We also employ the data augmentation technique that uses various permutations of elements in the quadruplet (Hu et al., 2022). The element permutations are ranked based on the prediction entropy of pre-trained T5-Base, and we use $P$ different permutations with minimal entropy for each quadruplet. If there exist multiple quadruplets for a single sentence, we concatenate them using a special symbol [SSEP] (2021a).
@@ D.2 Implementation Details
I am performing the ASQP task, which is the Subtask of ABSA. From now on, if I give you a sentence and a quadruplet, create a Reasoning for it. When creating, create to satisfy all of the following conditions: When proceeding with inference, extract the aspect term and the option term first and infer the aspect category and sentimental polarity based on them. When extracting aspect term and option term, an aspect category is judged by a combination of aspect term and opinion term, and sentimental polarity is judged by comprehensively considering everything. And don’t mention each element first, explain the reason first, and then create a rationale that mentions the element. At this time, do not number each element, but configure it to naturally lead to one paragraph. But if there are more than two quadruplets, please organize the description for each quadruplet. And please create a detailed description of each element in the composition of rationale. From now on, I’ll give you sentences and quadruplet sets as input.
Here are possible aspect category set: [’food prices’, ’food style_options’, ’service general’, ’drinks prices’, ’ambience general’, ’drinks quality’, ’location general’, ’restaurant prices’, ’restaurant general’, ’drinks style_options’, ’food general’, ’restaurant miscellaneous’, ’food quality’].
Here are possible sentiment polarity set: [’positive’,’negative’,’neutral’].

### v2

# Self-Consistent Reasoning-based Aspect Sentiment Quad Prediction with Extract-Then-Assign Strategy

@@ Abstract
In the task of aspect sentiment quad prediction (ASQP), generative methods for predicting sentiment quads have shown promising results. However, they still suffer from imprecise predictions and limited interpretability, caused by data scarcity and inadequate modeling of the quadruplet composition process. In this paper, we propose Self-Consistent Reasoning-based Aspect sentiment quadruple Prediction (SCRAP), optimizing its model to generate reasonings and the corresponding sentiment quadruplets in sequence. SCRAP adopts the Extract-Then-Assign reasoning strategy, which closely mimics human cognition. In the end, SCRAP significantly improves the model’s ability to handle complex reasoning tasks and correctly predict quadruplets through consistency voting, resulting in enhanced interpretability and accuracy in ASQP. [^1] Codes and datasets are available at https://github.com/jieyong99/SCRAP
@@ 1 Introduction
Aspect-based sentiment analysis (ABSA) refers to the task of identifying entity aspects and their associated sentiments (Pontiki et al., 2014). Among various ABSA tasks, the challenging task of predicting quadruplets, including aspect sentiment quad prediction (ASQP) (Zhang et al., 2021a) and aspect-category-opinion-sentiment (ACOS) (Cai et al., 2021), has garnered significant interest in that it can offer comprehensive aspect-level analysis. Specifically, a quadruplet consists of four sentiment elements: aspect term (at), opinion term (ot), aspect category (ac), and sentiment polarity (sp). Recent studies have developed powerful generative methods by fine-tuning language models (LMs) to sequentially generate sentiment quads (Zhang et al., 2021a; Hu et al., 2022; Gou et al., 2023).
@@
To tackle this challenge in ASQP, our approach introduces a two-step reasoning strategy, termed Extract-Then-Assign. This reasoning strategy initially extracts all aspect terms and opinion terms from an input sentence; subsequently, it assigns each aspect-opinion pair to both an aspect category and a sentiment polarity, utilizing predefined sets of categories and polarities. In Figure 1 Right, the process begins with a human extracting the information (at, ot) that is directly identifiable from the review sentence. Following this, the individual infers ac and sp based on the initially extracted components. We explicitly model these two steps of ASQP reasoning, which can account for the generation of complete quadruplets. This reasoning process, which closely mimics human cognition, allows us to enhance the accuracy and interpretability of quad prediction.
In this work, we propose a novel framework for Self-Consistent Reasoning-based Aspect-sentiment quad Prediction, named SCRAP, which predicts aspect sentiment quads based on the Extract-Then-Assign reasoning (Figure 2). The key idea is to distill the plausible reasoning ability from large language models (LLMs) into our ASQP model. To this end, we first collect diverse reasoning paths via Chain-of-Thought (CoT) prompting on LLMs Wei et al. (2022), and then optimize our model to generate the reasoning followed by quad prediction. Furthermore, SCRAP aggregates its diverse reasoning outputs based on the self-consistency (Wang et al., 2022). This allows for filtering out noisy outputs and achieving more accurate quad predictions.
@@ 2.1 Problem Formulation
Given an input sentence, aspect sentiment quad prediction (ASQP) aims to predict all aspect sentiment quads $\{(at,ot,ac,sp)\}$. The aspect term $at$ and opinion term $ot$ are detected within the sentence, while the aspect category $ac$ and sentiment polarity $sp$ are classified within their respective predefined sets. Please refer to Appendix A for the details.
@@ 2.2 ASQP Reasoning Generation
Mimicking the human cognition process, which first identifies terms and then infers their relations and semantics, we devise an Extract-Then-Assign reasoning strategy for ASQP. From an input sentence, it first extracts at and ot pairs, and subsequently infers the corresponding ac and sp by assigning them to elements within predefined sets of categories and polarities.
@@
We generate diverse reasoning paths using LLM based on the proposed reasoning strategy. We employ a few-shot Chain-of-Thought (CoT) prompting (Wei et al., 2022). Formally, given a sentence and its quadruplets, we generate $N$ reasoning paths $\mathcal{R}=[r_{1},r_{2},...,r_{N}]$ that explain how to reach the quadruplets from the sentence. Our prompt is designed to induce the Extract-Then-Assign process, facilitating the generation of plausible reasoning from the LLM. By leveraging the generated reasoning, we seek to provide rationales that guide and explain the ASQP task to our model, enhancing both accuracy and interpretability. The prompt can be found in the Appendix B.
Methods Rest15 Rest16
Pre Rec F1 Pre Rec F1
@@
[…]
SCRAP (Ours) 55.45 45.41 49.93 69.59 56.70 62.48
Table 1: ASQP performance comparison. Backbone model: T5-Base. The best and second-best results are in bold and underlined, respectively. $\dagger$ indicates the results reported from their original papers.
@@ 3.1 Experimental Settings
Dataset and evaluation metrics
We use two datasets, i.e., Rest15 and Rest16, widely used for the ASQP task (Zhang et al., 2021a). For reasoning generation (Sec.2.3), we use ChatGPT (gpt-3.5-turbo-16k). [^2] https://chat.openai.com/ As the evaluation metric, we mainly employ the F1 score with precision (Pre) and recall (Rec). A predicted quad is considered as correct if and only if its all elements are exactly the same as the ground-truth ones.
@@
We compare SCRAP with two discriminative methods, i.e., TAS-BERT Wan et al. (2020) and Extract-Classify Cai et al. (2021), as well as four competitive generative methods, i.e., GAS Zhang et al. (2021b), Paraphrase Zhang et al. (2021a), DLO Hu et al. (2022), and MvP Gou et al. (2023). For generative methods, we adopt T5-Base and T5-3B as backbone models. For SCRAP, we set $N=16$ and $P=5$ in common, $K=20$ (T5-Base) and $K=25$ (T5-3B) for Rest15, $K=15$ for Rest16. Refer to Appendix D for implementation details.
@@ 3.2 Results and Discussion
Table 1 and Figure 3 present the ASQP performance of various methods. Overall, SCRAP achieves competitive performance, and outperforms the baseline methods for large backbone models. Specifically, in Figure 3, SCRAP shows higher effectiveness when applied to a larger model, exhibiting the largest performance gap between T5-Base and T5-3B. A larger model generally has higher reasoning capability (Li et al., 2023), making it more suitable for SCRAP which leverages reasoning for ASQP. Furthermore, SCRAP generally exhibits high precision, as it filters out inconsistent predictions through consistency voting. When the task is performed with the T5-Base on the Rest15 dataset, the performance is relatively low due to the small dataset size and the use of a model with relatively poor inference capabilities. Lastly, we analyze the prediction errors of SCRAP and the best competitor, MvP. The prediction error is calculated for each component as (number of incorrect predictions) / (total number of predicted quads). In Figure 4 Left, the overall error rate of SCRAP is notably lower than MvP.
@@
We investigate the performance of SCRAP with varying $N$ and $K$, which control the number of reasoning paths used for the training and inference, respectively. In Figure 5, we observe that SCRAP generally achieves higher performance with a greater number of reasoning paths with respect to both $N$ and $K$, and the best performance is achieved by leveraging multiple paths for both training and inference. These results show that diverse reasoning for prediction is indeed beneficial in more accurate quad predictions. Nevertheless, if $K$ grows excessively, a considerable quantity of incomplete inferences are produced, leading to adverse effects on quad predictions and consequent performance deterioration. Therefore, it is crucial to determine the optimal value of multiple paths $K$ for inference.
@@
Figure 4 Mid and Right provide the case study. In Example 1, the model predicts the correct quad, and it is possible to interpret the process through which the correct answer was reached. Moreover, even if the model fails to predict the correct quad, we can still understand how the prediction failed based on the Extract-Then-Assign reasoning process. In Example 2, the model fails to extract the correct opinion term, and this leads to inaccurate polarity prediction. This interpretability helps to better understand the model behavior, which is an important strength of SCRAP. It demonstrates that the Extract-Then-Assign reasoning strategy is simple yet effective within the SCRAP framework.
@@ 4 Related Work
Many existing studies have focused on discriminative methods. Early works tried jointly detecting target-aspect-sentiment (Wan et al., 2020) or conducting ACOS with two-stage pipelines (Cai et al., 2020). Beginning with Zhang et al. (2021b), recent studies have started using generative methods. Zhang et al. (2021a) transforms outputs into natural language sentences, Hu et al. (2022) introduce data augmentation that uses various permutations of quad elements, and Gou et al. (2023) align training and inference with multi-view prompting.
@@
CoT prompting has shown high effectiveness in inducing models to generate reasoning before reaching an answer Wei et al. (2022); Kojima et al. (2022); Wang et al. (2022). Recent work has focused on distilling the reasoning capabilities of LLMs to smaller LMs Ho et al. (2022); Magister et al. (2022); Hsieh et al. (2023); they elicit rationales for the predictions from LLMs and utilize them to train the LMs, effectively improving their performance.
@@ 5 Conclusion
This paper aims to enhance quad prediction in ASQP using Extract-Then-Assign, which is the two-step reasoning strategy. To this end, we propose a SCRAP framework, which generates diverse predictions utilizing the Extract-Then-Assign and selects the final answer by filtering the inconsistent answers through consistency voting. Our framework is the first method to integrate reasoning into the ABSA task, not only achieving state-of-the-art performance, but also significantly enhancing the interpretability of the outputs produced by the model. This confirms its efficacy in predicting quadruplets through reasoning.
@@ 6 Limitations
Despite achieving state-of-the-art performance, our study has three limitations. Firstly, our current reasoning structure, designed to mimic human cognition, may not be the most advanced or optimal for the ASQP task. There could exist a more sophisticated reasoning structure that further enhances the task performance. Secondly, the effectiveness of our approach is affected by the size of the model to some extent. With small models having limited reasoning capabilities, our method may not exhibit satisfactory performance. Lastly, our study incurs higher computational costs for training and inference compared to previous studies, as it additionally leverages reasoning.
@@ 7 Ethical Statment
We utilize datasets that are widely recognized and previously employed in the scientific community, maintaining transparency and integrity in our experiments. Our methodologies and findings do not inflict harm upon any individuals or groups. We are cognizant of the potential biases in sentiment polarity predictions arising from the use of large pre-trained language models, as these models may mirror existing societal biases found in their training corpora (Tan and Celis, 2019). We acknowledge the importance of ongoing efforts to mitigate such biases. Furthermore, we underscore the necessity of continuous monitoring and evaluation to ensure that our smaller downstream models do not replicate or amplify the biases inherent in their larger language model counterparts.
Acknowledgements
This work was supported by the IITP grant funded by the Korea government (MSIT) (No. RS-2020-II201361) and the NRF grant funded by the Korea government (MSIT) (No. RS-2023-00244689).
@@ Appendix C Target Construction Details
We construct the prediction targets by combining the generated reasoning with the ground-truth quadruplets. To represent the quadruplets, we use special markers: [AT], [OT], [AC], [SP], which respectively denote at, ot, ac, sp, as done in (Hu et al., 2022). We also employ the data augmentation technique that uses various permutations of elements in the quadruplet (Hu et al., 2022). The element permutations are ranked based on the prediction entropy of pre-trained T5-Base, and we use $P$ different permutations with minimal entropy for each quadruplet. If there are multiple quadruplets for a single sentence, we concatenate them using a special symbol [SSEP] (2021a).
@@ D.2 Implementation Details
I am performing the ASQP task, which is the Subtask of ABSA. From now on, if I give you a sentence and a quadruplet, create a Reasoning for it. When creating, create to satisfy all of the following conditions: When proceeding with inference, extract the aspect term and the option term first and infer the aspect category and sentimental polarity based on them. When extracting aspect term and option term, an aspect category is judged by a combination of aspect term and opinion term, and sentimental polarity is judged by comprehensively considering everything. And don’t mention each element first, explain the reason first, and then create a rationale that mentions the element. At this time, do not number each element, but configure it to naturally lead to one paragraph. But if there are more than two quadruplets, please organize the description for each quadruplet. And please create a detailed description of each element in the composition of rationale. From now on, I’ll give you sentences and quadruplet sets as input. Here are possible aspect category set: [’food prices’, ’food style_options’, ’service general’, ’drinks prices’, ’ambience general’, ’drinks quality’, ’location general’, ’restaurant prices’, ’restaurant general’, ’drinks style_options’, ’food general’, ’restaurant miscellaneous’, ’food quality’]. Here are possible sentiment polarity set: [’positive’,’negative’,’neutral’].